(3Square root of 2 -9)(3 square root of 2+9) HOW DO I SOLVE THAT
Answer by richwmiller(17219)

((3Sqrt2) -9)((3Sqrt 2)+9)=
use FOIL
3Sqrt2*3Sqrt2=9*2=18
-9*3Sqrt+
9*3Sqrt2
-81
=
-63